OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. m = -4, point (2, -5

OpenStudy (anonymous):

y+5=-4(x-2) y+5=-4x+8 4x+y-3=0

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

I didn't get u....

OpenStudy (anonymous):

which one would be the equation?